How many liters of solution are needed to make a 1.66 M solution containing 333.44 grams of KMnO4?


1
Expert's answer
2021-04-09T04:05:09-0400

C=n/V

V=n/C

n=m/M

M=39+55+4*16=158(g/mol)

n=333.44/158=2.11(mol)

V=2.11/1.66=1.27(L)